OpenStudy (anonymous):

how do you find the radius of x^2+y^2+4x+6y+9=0. what is the working out to find the radius please ?

terenzreignz (terenzreignz):

You have to manipulate it to look like this... \[\Large (x-h)^2 + (y-k)^2 = r^2\] Then the radius would be r.

OpenStudy (anonymous):

yeah i know but the problem is that i keep making a mistake that i cant find. i end up with the right centre co-ordinates but the radius is wrong could you please do the working out so i can understand

terenzreignz (terenzreignz):

Of course. It all boils down to a nifty little process called "completing the square" Let's write your equation here... \[\Large \color{red}{x^2}+\color{blue}{y^2}+ \color{red}{4x}+\color{blue}{6y}+\color{black}9 = 0\]

terenzreignz (terenzreignz):

You'll want to group the x's and the y's together, like so... \[\Large \color{red}{x^2+4x}+\color{blue}{y^2+6y}+9 = 0\]

terenzreignz (terenzreignz):

Now, to complete the square given the form \[\Large z^2 + pz\] You take the coefficient of the term where the variable is raised to the exponent 1, and that's p. Get half of it, and then square it. That's \(\Large \left(\frac{p}{2}\right)^2\) You add it to the expression, but ALSO SUBTRACT IT to keep it equal to the original expression.... \[\Large = z^2 + pz \color{green}{+\frac{p^2}{4}}\color{orange}{-\frac{p^2}{4}}\] Such that it is now equal to... \[\Large \left(z+\frac{p}2\right)^2 -\frac{p^2}4\]

terenzreignz (terenzreignz):

So, let's complete the square, shall we? I'll do the x, you do the y. \[\Large \color{red}{x^2+4x}+\color{blue}{y^2+6y}+9 = 0\] So, the coefficient of the x (with no exponent) is 4. Half of 4 is 2, and square of 2 is 4. So we add 4 and we also add 4 to the other side, to balance it out... \[\Large \color{red}{x^2+4x}\color{orange}{+4}+\color{blue}{y^2+6y}+9 = 0\color{orange}{+4}\]

OpenStudy (anonymous):

I cant solve the part where i got upto which is (x+2)^2 + (y+3)^2=4+9+9

terenzreignz (terenzreignz):

Ahh... I see :) why do you have two "+9"s there?

terenzreignz (terenzreignz):

One of those 9's came from the left side, I'm sure of it. It was +9 on the left side, so doesn't that mean it becomes -9 when you bring it to the right? ;)

terenzreignz (terenzreignz):

Simplifying... \[\Large \color{red}{(x+2)^2}+\color{blue}{y^2+6y}+9 = 4\] Now, completing the square of the y's \[\Large \color{red}{(x+2)^2}+\color{blue}{y^2+6y}\color{green}{+9}+9 = 4\color{green}{+9}\] Simplifying again... \[\Large \color{red}{(x+2)^2}+\color{blue}{(y+3)^2}+9 = 13\]

terenzreignz (terenzreignz):

Still in doubt? :)

OpenStudy (anonymous):

thank you very much you've helped me alot :)

terenzreignz (terenzreignz):

No problem :)

OpenStudy (amistre64):

\[\Large \color{red}{x^2+4x}+\color{blue}{y^2+6y+9} = 0\] \[\Large \color{red}{x^2+4x+4}+\color{blue}{y^2+6y+9} = 4\]

OpenStudy (amistre64):

the y parts are already a complete square

terenzreignz (terenzreignz):

Me and my formulaic self :3